A pharmacist has 6L of a mixture that is 20% alcohol. How much pure alcohol must be added to obtain a mixture that is 70% alcohol?
Answer by ankor@dixie-net.com(22740)

A pharmacist has 6L of a mixture that is 20% alcohol.
How much pure alcohol must be added to obtain a mixture that is 70% alcohol?
:
let x = amt of pure alcohol required
then
(6-x) = amt of 20% alcohol required
:
.20(6-x) + x = .70(6)
1.2 - .2x + x = 4.2
.8x = 4.2 - 1.2
.8x = 3
x = 3/.8
x = 3.75 L of pure alcohol required
:
:
Check: .2(2.25) + 3.75 = .7(6)
.45 + 3.75 = 4.2